One man starts out from a city and walks North at the rate of 3 kilometers per hour. Three hours later, a second man starts from the same place and walks the same route at 4 kilometers per hour. How far will he have to go before he overtakes the first man?
Answer by greenestamps(13198)



The first man walks 3 km/h for 3 hours before the second man starts, traveling a distance of 3*3=9km.

The rate at which the second man overtakes the first man is the difference in their walking rates, which is 1km/h.

To make up a distance of 9km at a rate of 1km/h takes 9/1=9 hours.

In 9 hours the second man walks 9*4=36km. (And the first man by then has walked 12*3=36km.)

ANSWER: The second man walks 36km before he overtakes the first man.